# FreqPolicy: Frequency Autoregressive Visuomotor Policy with Continuous Tokens

### v1

@@ F.2 Discussion on Frequency Domain
In the preceding discussions, we have preliminarily shown that the action signals in robotic manipulation tasks exhibit significant compressibility in the frequency domain, with most critical information concentrated in the lower frequency bands. This section aims to provide a more in-depth discussion and analysis of the frequency-domain characteristics of action signals, based on a broader set of tasks and more detailed visualizations (as shown in Figures 10 to 30). These supplementary figures provide action visualizations (a), frequency band energy heatmaps for each action dimension (b), and success rate curves for actions reconstructed with varying frequency ratios (c) for each task, thereby offering more robust support for our core arguments.

### v2

@@ F.2 Discussion on Frequency Domain
In the preceding discussions, we have preliminarily shown that the action signals in robotic manipulation tasks exhibit significant compressibility in the frequency domain, with most critical information concentrated in the lower frequency bands. To make this statement precise, we quantify “compressibility” by the energy proportion:
$E(p)=\sum\nolimits_{k=0}^{\lfloor(N-1)p\%\rfloor}|X_{k}|^{2}\Big/\sum\nolimits_{k=0}^{N-1}|X_{k}|^{2},$ (4)
i.e. the share of total signal energy that is preserved when only the lowest $p\%$ DCT coefficients are retained (see Figure 1).
This section aims to provide a more in-depth discussion and analysis of the frequency-domain characteristics of action signals, based on a broader set of tasks and more detailed visualizations (as shown in Figures 10 to 30). These supplementary figures provide action visualizations (a), frequency band energy heatmaps for each action dimension (b), and success rate curves for actions reconstructed with varying frequency ratios (c) for each task, thereby offering more robust support for our core arguments.